# Bayesian Optimization for Controlled Image Editing via LLMs

### v1

@@ 1 Introduction
Controllable synthesis in generation technology Hertz et al. (2022); Brooks et al. (2023); Patashnik et al. (2021) has recently attracted significant attention due to its expanded range of applications. Models such as Pix2Pix and CycleGAN have demonstrated the ability to transform images from one domain to another, effectively applying controllable synthesis to tasks like style transfer and image enhancement Isola et al. (2017); Zhu et al. (2017). Recent advances like ZONE have enabled instruction-driven modifications without pre-defined training samples Li et al. (2024a), while new frameworks have emerged for intuitive and localized image editing by manipulating internal attention mechanisms Brooks et al. (2023). This represents a shift towards more granular control over AI-generated content, enabling precise, region-specific alterations without additional input masks.
However, despite these advancements, existing methods face several critical challenges. First, most state-of-the-art local editing methods heavily rely on mask priors to constrain the editing regions—either through manual input or derived from attention map analysis and semantic segmentation—making them less accessible for non-expert users. Additionally, methods like cross-attention control and diffusion models often encounter challenges in fine-tuning model parameters to align with user requirements, resulting in a disparity between desired and actual outputs. These issues are particularly pronounced in applications that demand detailed modifications based on user instructions.
@@ 1.1 Related Work
The Instruct Pix2Pix framework represents a significant advancement in the field of image editing Brooks et al. (2023). This model has been widely used in various image-to-image generation tasks, such as converting hand-drawn sketches into photographs M Shetty K Raghavendra (2022), transforming abstract maps into realistic map images Li et al. (2024b) and de-noising images taken in harsh environments for crowd counting Khan et al. (2023). Instruct Pix2Pix employs Classifier-Free Guidance (CFG) for both image and text conditions, adjusting the weights of these inputs to control the generated output. It enables users to use natural language instructions for image editing, leveraging the model’s ability to implement detailed modifications. The system manipulates the internal attention mechanisms of generative models, offering precise alterations without the need for additional input masks. Innovations such as DALLE-3 and CLIP integrate multi-modal learning, leveraging large-scale text and image datasets to enhance contextual understanding and generation capabilities Betker et al. (2023); Radford et al. (2021).
However, a common limitation persists across most state-of-the-art approaches: they typically rely on some form of mask guidance to achieve precise local editing. Whether through manual mask annotations Hertz et al. (2022), attention map analysis Li et al. (2024a), or semantic segmentation, the dependence on mask priors creates a barrier for non-expert users and limits the flexibility of these systems. BayesGenie addresses this limitation by enhancing the diffusion model with automated parameter optimization, enabling it to follow instructions more accurately while preserving the high quality of the generated images without requiring any form of mask guidance.
@@ 3.2 Experimental Setup
Most state-of-the-art local editing methods, including ZONE Li et al. (2024a), heavily rely on mask priors to constrain the editing regions—either through manual input or derived from operations such as attention map analysis and semantic segmentation. In contrast, our method, to the best of our knowledge, is the first to aim for localized editing without relying on any form of mask guidance. As such, comparing our approach with mask-guided methods would neither be fair nor meaningful.
@@ 5 Conclusion
6 Limitations
While our proposed method, BayesGenie, demonstrates substantial improvements in controllable image editing, it is not without limitations. First, the reliance on large language models (LLMs) like GPT-4 and Claude 3.5 introduces potential biases and limitations inherent to these models. Specifically, LLMs can exhibit hallucination effects, particularly when evaluating subtle image modifications, leading to inconsistencies in scoring or inappropriate recommendations for image adjustments, which could impact the overall precision of the edited image. Future directions could include fine-tuning open-source LLMs, such as Llava, to develop a more stable and consistent model for both image modification evaluation and textual prompt generation. By improving the alignment between the generated images and user intentions, a fine-tuned LLM could provide a more reliable baseline for both assessing and generating image outputs, thus enhancing the overall effectiveness of the system.

### v2

@@ 1 Introduction
Controllable synthesis in generation technology Guan et al. (2025); Yao et al. (2024); Hertz et al. (2022); Brooks et al. (2023); Patashnik et al. (2021); Jiang et al. (2024) has recently attracted significant attention due to its expanded range of applications. Models such as Pix2Pix and CycleGAN have demonstrated the ability to transform images from one domain to another, effectively applying controllable synthesis to tasks like style transfer and image enhancement Isola et al. (2017); Zhu et al. (2017). Recent advances like ZONE have enabled instruction-driven modifications without pre-defined training samples Li et al. (2024b), while new frameworks have emerged for intuitive and localized image editing by manipulating internal attention mechanisms Brooks et al. (2023). This represents a shift towards more granular control over AI-generated content, enabling precise, region-specific alterations without additional input masks.
However, despite these advancements, existing methods face several critical challenges. First, most state-of-the-art local editing methods heavily rely on mask priors to constrain the editing regions—either through manual input or derived from attention map analysis and semantic segmentation—making them less accessible for non-expert users. Additionally, methods like cross-attention control and diffusion models often encounter challenges in fine-tuning model parameters to align with user requirements, resulting in a disparity between desired and actual outputs. These issues are particularly pronounced in applications that demand detailed modifications based on user instructions. With the LLMs-Driven adaptation, recent various applications Cai et al. (2024b, a); Li et al. (2024a); Shi et al. (2025); Liu et al. (2024); Shi et al. (2024) have advanced developed.
@@ 1.1 Related Work
The Instruct Pix2Pix framework represents a significant advancement in the field of image editing Brooks et al. (2023). This model has been widely used in various image-to-image generation tasks, such as converting hand-drawn sketches into photographs M Shetty K Raghavendra (2022), transforming abstract maps into realistic map images Li et al. (2024c) and de-noising images taken in harsh environments for crowd counting Khan et al. (2023). Instruct Pix2Pix employs Classifier-Free Guidance (CFG) for both image and text conditions, adjusting the weights of these inputs to control the generated output. It enables users to use natural language instructions for image editing, leveraging the model’s ability to implement detailed modifications. The system manipulates the internal attention mechanisms of generative models, offering precise alterations without the need for additional input masks. Innovations such as DALLE-3 and CLIP integrate multi-modal learning, leveraging large-scale text and image datasets to enhance contextual understanding and generation capabilities Betker et al. (2023); Radford et al. (2021).
However, a common limitation persists across most state-of-the-art approaches: they typically rely on some form of mask guidance to achieve precise local editing. Whether through manual mask annotations Hertz et al. (2022), attention map analysis Li et al. (2024b), or semantic segmentation, the dependence on mask priors creates a barrier for non-expert users and limits the flexibility of these systems. BayesGenie addresses this limitation by enhancing the diffusion model with automated parameter optimization, enabling it to follow instructions more accurately while preserving the high quality of the generated images without requiring any form of mask guidance.
@@ 3.2 Experimental Setup
Most state-of-the-art local editing methods, including ZONE Li et al. (2024b), heavily rely on mask priors to constrain the editing regions—either through manual input or derived from operations such as attention map analysis and semantic segmentation. In contrast, our method, to the best of our knowledge, is the first to aim for localized editing without relying on any form of mask guidance. As such, comparing our approach with mask-guided methods would neither be fair nor meaningful.